When you place a heavy encyclopedia, of mass 8 kg, on a kitchen scale (a spring scale)
you notice that before coming to equilibrium, the pointer of the scale oscillates back and forth around
the equilibrium position a few times with a period of 0.4 s. What is the effective spring constant of the
internal spring of the kitchen scale? (Neglect other masses in the scale.)
1
Expert's answer
2020-03-13T11:06:48-0400

"T=2\pi\sqrt{\frac{m}{k}} \to k=\frac{4\pi^2m}{T^2}=\frac{4\cdot3.14^2\cdot8}{0.4^2}=1974\approx2000" "N\/m"